6. The OECD (Organization for Economic Cooperation and Development) consists of 20
wealthy countries. The prediction equation relating y=child poverty rate to x=social expenditure
(as a percent of GDP) is yhat=22-1.3x.
a. Interpret the slope.
b. Find the predicted poverty rates for the U.S. and Denmark. The x-value for U.S. is
2% and for Denmark the x-value is 16%.
Answer #1

a)

slope tells us that for 1% increase in social expenditure ( (as a percent of GDP ) ; on average child poverty reduce by 1.3 units

b)

predicted poverty rates for the U.S =22-1.3*2 =19.4

predicted poverty rates for the Denmark =22-1.3*16 =1.2